please help me solve: F(x)=3x g(x)=x+7 solve f(2)-g(3)
Is it: f(2.3x)-g(3.x+7)= 6x-3x+21
3x+21divided by 3 = x+7
The . is a times sign!

Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

please help me solve:
---
It's not solving. Equations are solved, this is not one.
------------
Use * (Shift 8) for multiply, or the "times sign."
--------------
F(x)=3x g(x)=x+7 solve f(2)-g(3)
Assuming f(x) and F(x) are the same:
---
f(x) = 3x
To find f(2), sub 2 for each x
f(2) = 3*2 = 6
---
g(x) = x+7
Wherever there's an x, sub a 3
g(3) = 3+7 = 10
----
===================
f(2) - g(3) = 6-10 = -4
----
You're making it more complicated than needed. Divided? Calm down.
--------------------
Is it: f(2.3x)-g(3.x+7)= 6x-3x+21
3x+21divided by 3 = x+7
The . is a times sign!

Answer by jim_thompson5910(35256)

First compute f(2)

f(x) = 3x

f(2) = 3*2 ... replace EVERY x with 2

f(2) = 6

-------------------------------------------------------

Now compute g(3)

g(x) = x+7

g(3) = 3+7 ... replace EVERY x with 3

g(3) = 10

-------------------------------------------------------

So,

f(2) - g(3) = 6 - 10 = -4

In short, f(2) - g(3) = -4


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim